My sister wanted help with her Algebra but I haven't had any algebra in almost 10 years. I have just enough knowledge for it to be dangerous :) Thanks
5x-(2x-10)=35

Answer by sprolden(40)

5x-(2x-10)=35
1. Distribute the (-) in front of the parenthesis by multiplying -1 times everything in the parenthesis. 5x-2x+10=35
2. Subtract the x terms. You'll get 3x+10=35
3. Subtract 10 from both sides. 3x = 25
4. Divide by 7 on both sides. x = 25/3 OR 8.33333333